OpenStudy (anonymous):

y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (anonymous):

3(x-1)=x 3+1=y so then when you add or subtract you need to put the equations together and then solve

OpenStudy (anonymous):

3(x-1)=x 3+1=y so then when you add or subtract you need to put the equations together and then solve

OpenStudy (campbell_st):

both equations are equal to y, so replace y in the 1st equation with the right hand side of the 2nd equation so you get x - 1 = 3x + 3 now you can solve for x. when you get x, substitute it into either of the original equations to find y. hope it helps

OpenStudy (campbell_st):

that's the substitution method, all you are doing is finding where the 2 lines intersect. on the number plane.

OpenStudy (anonymous):

Yeah, there are 2 ways to solve a system of equations. :)